User name Carlos
Log entry time 00:20:31 on October15,2004
Entry number 134210
keyword=Swing shift summary
16:00 Control access terminated
16:00 Beam studies (no beam in the hall)
22:00 MCC called. Updated estimate beam in 1h.
22:55 MCC called. They'll try to send some pulsed beam to the hall and later CW. We moved our target from carbon (single foil) to empty.
23:55 MMC sending CW beam to the Hall. Performed harp scan (see Halog entry) and tuned BPMs.